3. Determine iz(t) for p0 for the circuit shown. 20 mH L 50 ms os
Answer #1

® Given cércuit 2o mnt mm M 50ms 342 We need to consider the three time intervals tso, ostssoms and t> 50ml respectively. ForFor osts 5oms Switch s, is closed. - Inom - 2014 { 402 no current flows in 422 since aurent always choose loro resistance patFor t>soms Switch, s, is closed m m it) This sudden change does not affect the inductor current because the current cannot chHence (t) = 160) + [{soms)-i cey] e{t-som) Czce) = 12 + (1124-12] Ct-sex8).ous 2,67= 12 +0.798-400 , tz soms